OpenStudy (anonymous):

If a,b are the roots of ax^2+bx+c=0, find the value of (a^3+b^3)-3ab(a^2+b^2) . . ANSWER = (3abc+6ac^2-b^3-3b^2c)/a^3

OpenStudy (harsimran_hs4):

you have a + b = -b/a and ab = c/a and a^3 + b^3 = (a+b)^3 - 3ab(a+b) a^2 + b^2 = (a+b)^2 - 2ab can you proceed now?

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

(a+b)^3 - 3ab(a+b) - 3ab ((a+b)^2 - 2ab)

OpenStudy (anonymous):

this will be the whole expansion right?

OpenStudy (harsimran_hs4):

yes!!

OpenStudy (anonymous):

I am stuck here

OpenStudy (anonymous):

\[-\frac{ b^3 }{ a^3 }+\frac{ 3bc }{ a^2 }-\frac{ 3c }{ a }(\frac{ b^2-2ac }{ a^2 })\]

OpenStudy (anonymous):

ok I've done it